### 25 Apr 2026 v1

# Boundedness of commutator generated by fractional integral operator and Orlicz-BMO function

Abstract.
For $\alpha\in(0,n)$ and a growth function $\varphi:[0,\infty)\rightarrow[0,\infty)$, it is proved that the commutator $b,I_{\alpha}]$ generated by fractional integral operator $I_{\alpha}$ and Orlicz $\mathrm{BMO}$ function $b$ is bounded from Orlicz-Hardy space $H_{b}^{\varphi}(\mathbb{R}^{n})$ to Lebesgue space $L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})$, where $H_{b}^{\varphi}(\mathbb{R}^{n})$ is a suitable Orlicz-Hardy space. Moreover, the authors also establish that the boundedness of commutator $b,I_{\alpha}]$ from Orlicz-Hardy space $H^{\varphi}(\mathbb{R}^{n})$ to weak Lebesgue space $L^{\frac{n}{n-\alpha},\infty}(\mathbb{R}^{n})$.
Key words and phrases:
Fractional integral operator; Commutator; Orlicz-Hardy space; Orlicz-$\mathrm{BMO}$ space; Boundedness.
2020 Mathematics Subject Classification:
31C40; 42B30; 42B35.
Chenglong Fang is supported by the National Natural Science Foundation of China (# 12501119) and the Fujian Provincial Natural Science Foundation Young Scientist Innovation Project (# 2025J08035).
Corresponding author
1. Introduction
For $0<\alpha<n$, the fractional integral operator $I_{\alpha}$ is defined as follows:
$I_{\alpha}(g)(x):=\int_{\mathbb{R}^{n}}\frac{g(z)}{\left|x-z\right|^{n-\alpha}}\,dz\quad\text{for all}\,x\in{{{\mathbb{R}}}^{n}}.$
Two classical results (see [9, Theorem 1.2.3] and [11, Theorem 4]) asserted that there exists constants $C$ such that
(1.1) $\|I_{\alpha}g\|_{L^{\frac{n}{n-\alpha},\,{\infty}}(\mathbb{R}^{n})}\leq C\|g\|_{L^{1}({{{\mathbb{R}}}^{n}})}$
and
$\|I_{\alpha}g\|_{L^{q}(\mathbb{R}^{n})}\leq C\|g\|_{L^{p}({{{\mathbb{R}}}^{n}})}$
with $1<p<q<\infty$ and $1/q-1/p=\alpha/n$. In the framework of Hardy spaces, the boundedness of $I_{\alpha}$ is considered by Stein and Weiss [21, Theorem G] in their pioneering work on multi-variable harmonic functions, they established that
(1.2) $\|I_{\alpha}g\|_{L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})}\leq C\|g\|_{H^{1}({{{\mathbb{R}}}^{n}})},$
which reveals the compatibility between the regularity of fractional integral operator and the atomic structure of Hardy spaces.
Given a locally integrable function $b$ and a sublinear operator $T$, the commutator is denoted by $[b,T]$, it is given by
[b,T]g:=bTgT(bg),[b,T]g:=bTg-T(bg),
where $g$ is a continuous function whose support is compact. A famous theorem of Chanillo [3] characterizes the $\mathrm{BMO}({{{\mathbb{R}}}^{n}})$ space via the boundedness of $b,I_{\alpha}]$, that is, $b\in\mathrm{BMO}({{{\mathbb{R}}}^{n}})$ if and only if the commutator $b,I_{\alpha}]$ is bounded from $L^{p}({{{\mathbb{R}}}^{n}})$ to $L^{q}({{{\mathbb{R}}}^{n}})$, where $0<\alpha<n$ and $1<p<\frac{n}{\alpha}$ with $1/q=1/p-\alpha/n$. Paluszyński [19, Theorem 1.1] proved that $b,I_{\alpha}]$ does not map from $H^{1}({{{\mathbb{R}}}^{n}})$ into $L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$ continuously. Later, Pérez [20] introduced a subspace $\mathcal{H}^{1}_{b}({{{\mathbb{R}}}^{n}})\subset H^{1}({{{\mathbb{R}}}^{n}})$ in the study of commutator of Calderón-Zygmund operator. Prompted by the bilinear splitting technique applied to the pair of Hardy space $H^{1}({{{\mathbb{R}}}^{n}})$ and $\mathrm{BMO}$ space, which is originally introduced by Bonami and Grellier [1], an important work is that Ky [13] found a suitable subspace $H_{b}^{1}({{{\mathbb{R}}}^{n}})$ of $H^{1}({{{\mathbb{R}}}^{n}})$ and established that $b,I_{\alpha}]$ have boundedness from $H_{b}^{1}({{{\mathbb{R}}}^{n}})$ to $L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$. Moreover, Ky [13] also proved that the $b,I_{\alpha}]$ maps $H^{1}({{{\mathbb{R}}}^{n}})$ continuously into the weak Lebesgue space $L^{\frac{n}{n-\alpha},\,\infty}({{{\mathbb{R}}}^{n}})$. For more related results, we refer to the interested reader to [3, 4, 10, 15, 20].
Motivated by the above works, the present paper studies the boundedness of commutator generated by Orlicz-$\mathrm{BMO}$ function and fractional integral operator within the framework of Orlicz-type spaces. To formulate our main results precisely, it is necessary to establish some definitions and notation.
A non-decreasing function $\varphi:[0,\infty)\rightarrow[0,\infty)$ is termed Orlicz function if it satisfies: $\varphi(0)=0$, $\varphi(s)>0$ for all $s>0$ and $lim_{s\rightarrow\infty}\varphi(s)=\infty$. For Orlicz function $\varphi$, it is of lower type $p_{1}$ (with $p_{1}\in[0,\infty)$) whenever one can find a constant $C_{p_{1}}>1$ such that
$\varphi(st)\leq C_{p_{1}}s^{p_{1}}\varphi(t)\qquad\forall\,t\geq 0,\;s\in(0,1].$
The symbol
$i(\varphi):=\sup\{p_{1}\geq 0:\varphi\,\text{is of lower type}\,p_{1}\}$
represents the critical lower type index of $\varphi$. For Orlicz function $\varphi$, it is of upper type $p_{2}$ (with $p_{2}\in[0,\infty)$) whenever one can find a constant $C_{p_{2}}>1$ such that
$\varphi(st)\leq C_{p_{2}}s^{p_{2}}\varphi(t)\qquad\forall\,t\geq 0,\,\;s\in[1,\infty).$
The symbol
$I(\varphi):=\inf\{p_{2}\geq 0:\varphi\,\text{is of upper type}\,p_{2}\}$
represents the critical upper type index of $\varphi$.
Definition 1.1.
We term $\varphi:[0,\infty)\to[0,\infty)$ a growth function if it is of upper type $1$ and lower type $p$ with $p\in(0,1]$.
Definition 1.2 ([14]).
For a given Orlicz function $\varphi$, the space $\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})$ stands for the set of all locally integrable functions $f$ on $\mathbb{R}^{n}$ satisfying
$\|g\|_{\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})}:=\sup_{B\subset\mathbb{R}^{n}}\frac{1}{\|\mathbf{1}_{B}\|_{L^{\varphi}(\mathbb{R}^{n})}}\int_{B}|g(x)-g_{B}|\,dx<\infty,$
where the supremum ranges over all balls $B$ on $\mathbb{R}^{n}$ and
$g_{B}:=\frac{1}{|B|}\int_{B}g(x)\,dx.$
If $\varphi(t)=t$, $\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})=\mathrm{BMO}(\mathbb{R}^{n})$. For more research on $\mathrm{BMO}$ type spaces, see [6, 7, 8, 22].
The Orlicz-Lebesgue space $L^{\varphi}({{{\mathbb{R}}}^{n}})$ contains all locally integrable functions $g$ for which
$\|g\|_{L^{\varphi}({{{\mathbb{R}}}^{n}})}:=\inf\left\{\lambda>0:\,\int_{{{{\mathbb{R}}}^{n}}}\varphi(|g(x)|/\lambda)\,dx\leq 1\right\}<\infty.$
We denote by $\mathcal{S}(\mathbb{R}^{n})$ the Schwartz class of rapidly decreasing smooth functions on $\mathbb{R}^{n}$, and by $\mathcal{S}^{\prime}(\mathbb{R}^{n})$ its topological dual (space of tempered distributions). Define
$\mathcal{S}_{0}(\mathbb{R}^{n}):=\left\{\psi\in\mathcal{S}(\mathbb{R}^{n}):\,\|\psi\|=\sup_{\begin{subarray}{c}x\in\mathbb{R}^{n}\\ |\alpha|\leq 1\end{subarray}}(1+|x|)^{2(n+1)}\,|\partial_{x}^{\alpha}\psi(x)|\leq 1\right\}.$
Let $g\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$ and $\psi_{t}(\cdot)=t^{-n}\psi(\cdot/t)$. The nontangential grand maximal function $\mathfrak{M}g$ is defined by
$\mathfrak{M}(g)(x):=\sup_{\psi\in S_{0}(\mathbb{R}^{n})}\;\sup_{\begin{subarray}{c}y\in\mathbb{R}^{n}\\ |y-x|<t\end{subarray}}|(g*\psi_{t})(y)|\quad\text{for all}\,x\in\mathbb{R}^{n}.$
Definition 1.3.
The Orlicz-Hardy space $H^{\varphi}(\mathbb{R}^{n})$ is the collection of all tempered distributions $g$ such that $\mathfrak{M}g$ belongs to $L^{\varphi}(\mathbb{R}^{n})$. Its (quasi-)norm is
$\|g\|_{H^{\varphi}}:=\|\mathfrak{M}g\|_{L^{\varphi}}.$
Definition 1.4.
Suppose that $\varphi$ is a growth function satisfying
$\frac{n}{n+1}<i(\varphi)\leq 1.$
Given $b\in\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})$, we denote by $H_{b}^{\varphi}({{{\mathbb{R}}}^{n}})$ the collection of all functions $g\in H^{\varphi}({{{\mathbb{R}}}^{n}})$ for which the function
$b,\mathfrak{M}]g:=\mathfrak{M}\left((b-b(\cdot))g(\cdot)\right)$
is an element of $L^{1}(\mathbb{R}^{n})$. For any $g\in H_{b}^{\varphi}({{{\mathbb{R}}}^{n}})$, define
(1.3) $\|g\|_{H_{b}^{\varphi}(\mathbb{R}^{n})}:=\|g\|_{H^{\varphi}(\mathbb{R}^{n})}\|b\|_{\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})}+\|[b,\mathfrak{M}]\|_{L^{1}(\mathbb{R}^{n})}.$
The subsequent primary results offer a generalisation of the preceding results on the boundedness in Ky [13] to the Orlicz setting.
Theorem 1.5.
Suppose that $0<\alpha<n$ and $\varphi$ is growth function with
(1.4) $n/(n+1)<i(\varphi)\leq I(\varphi)<1$
and
(1.5) $\begin{cases}1/I(\varphi)>1/i(\varphi)-1\quad&\mbox{if}\ n=1;\\ 1/I(\varphi)>(\lfloor n(1/i(\varphi)-1)\rfloor+n-1)/n\quad&\mbox{if}\ n\geq 2.\end{cases}$
If $b\in\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})$ and $g\in H_{b}^{\varphi}({{{\mathbb{R}}}^{n}})$, then there exists a positive constant $C$ such that
$\|[b,I_{\alpha}]g\|_{L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})}\leq C\|g\|_{H_{b}^{\varphi}({{{\mathbb{R}}}^{n}})}.$
Theorem 1.6.
Under the same assumptions on $\varphi$, $\alpha$, and $b$ as in Theorem 1.5, if $g\in H^{\varphi}({{{\mathbb{R}}}^{n}})$, then there exists constant $C>0$ such that
$\|[b,I_{\alpha}]g\|_{L^{\frac{n}{n-\alpha},\,\infty}({{{\mathbb{R}}}^{n}})}\leq C\|g\|_{H^{\varphi}({{{\mathbb{R}}}^{n}})}.$
Remark 1.7.
Theorems 1.5 and 1.6 are a further extension for the results of Ky [13, Corollary 8.1 and Theorem 8.2] in the Orlicz-type spaces framework. A study related to the Calderón-Zygmund operator, see [5].
In Section 2, we introduce the operator class $\mathcal{K}_{\varphi,\,\alpha}$ and prove $I_{\alpha}\in\mathcal{K}_{\varphi,\,\alpha}$. Then we establish the decomposition for commutator via wavelets in Section 3. Finally, Theorems 1.5 and 1.6 are verified in Section 4.
Finally, we give the notational rules adopted in this work. Let $\mathbb{N}:=\{1,2,\cdots\}$. For every $\in[0,\infty), t\lfloor t\rfloor$ denotes the supremum of all integers $k$ with $k\leq t$. For $x_{0}\in{{{\mathbb{R}}}^{n}}$ and $r\in(0,\infty)$, define $B(x_{0},r):=\{z\in{{{\mathbb{R}}}^{n}}:\,|z-x_{0}|<r\}$. The symbol $g_{1}\lesssim g_{2}$ (or $g_{2}\gtrsim g_{1})$ means $g_{1}\leq Cg_{2}$ (or $g_{2}\geq Cg_{1}$), and $g_{1}\approx g_{2}$ means both $g_{1}\lesssim g_{2}$ and $g_{2}\lesssim g_{1}$ hold.
2. Class $\mathcal{K}_{\varphi,\,\alpha}$ and typical example
In order to give the definition of $\mathcal{K}_{\varphi,\,\alpha}$, we start by recalling from [14] the atomic description of Orlicz-Hardy spaces. Subsequently, a typical example is presented in $\mathcal{K}_{\varphi,\,\alpha}$.
For ball $B\subset\mathbb{R}^{n}$ and $q\in[1,\infty]$, the symbol $L^{q}_{B}$ means the space of all measurable functions $g$ on $\mathbb{R}^{n}$, which $g$ satisfies $\operatorname{supp}g\subset B$, the norm of $L^{q}_{B}$ is defined by
(2.6) $\|g\|_{L^{q}_{B}}:=\begin{cases}\displaystyle\left(\frac{1}{|B|}\int_{B}|g(x)|^{q}\,dx\right)^{\!1/q}<\infty,&1\leq q<\infty;\\[10.76385pt] \displaystyle\operatorname{ess\,sup}_{x\in B}|g(x)|<\infty,&q=\infty.\end{cases}$
Definition 2.1.
Suppose that $\varphi$ is growth function with the critical lower type exponent $i(\varphi)$. A triple $(\varphi,q,s)$ is termed admissible if $q\in(1,\infty]$ and the positive integer $s$ satisfies
$s\geq m(\varphi):=\lfloor n(1/i(\varphi)-1)\rfloor.$
A locally integrable function $a$ on ${{\mathbb{R}}}^{n}$} is termed $(\varphi,q,s)$-atom if the next conditions are valid:
  1. (i)
    $\operatorname{supp}a\subset B$ for some ball $B\subset\mathbb{R}^{n}$;
  2. (ii)
    $\left\|a\right\|_{L^{q}_{B}}\leq\left\|{\mathbf{1}}_{B}\right\|^{-1}_{L^{\varphi}({{{\mathbb{R}}}^{n}})}$;
  3. (iii)
    $\int_{\mathbb{R}^{n}}x^{\gamma}a(x)\,dx=0$ for every multi-indices $\gamma\in{\mathbb{Z}}_{+}^{n}$ with $|\gamma|\leq s$.
Definition 2.2.
For $1\leq q\leq\infty$ and $s\geq m(\varphi):=\lfloor n(1/i(\varphi)-1)\rfloor$, suppose that $\{b_{j}\}_{j}$ is a sequence of multiples of $(\varphi,q,s)$-atom, each of which is supported in ball $B_{j}$. Define
$\Lambda_{q}\left(\{b_{j}\}\right):=\inf\left\{\lambda>0:\sum_{j}\varphi\left(\frac{\|b_{j}\|_{L_{B_{j}}^{q}}}{\lambda}\right)|B_{j}|\leq 1\right\}.$
Definition 2.3.
Let $0<\alpha<n$ and $\varphi$ be growth function such that $n/(n+1)<i(\varphi)\leq 1$. The $\mathcal{K}_{\varphi,\,\alpha}$ stands for the set consisting of all sublinear operator $R$ that fulfills the following three conditions below:
  1. (i)
    $R$ is bounded from $H^{1}(\mathbb{R}^{n})$ to $L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$;
  2. (ii)
    $R$ is bounded from $L^{1}(\mathbb{R}^{n})$ to $L^{\frac{n}{n-\alpha},\,\infty}({{{\mathbb{R}}}^{n}})$;
  3. (iii)
    For every $b\in\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})$ and every $(\varphi,q,0)$-atom $a$ whose support is contained in a ball $B\subset\mathbb{R}^{n}$ with $q\in(1,n/\alpha)$, there exists constant $C>0$ such that
    $\|(b-b_{B})Ra\|_{L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})}\leq C\|a\|_{L^{q}_{B}}\|{\mathbf{1}_{B}\|_{L^{\varphi}({{{\mathbb{R}}}^{n}})}\|b\|_{\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})}}.$
To facilitate subsequent proof, it is proposed that the result in [5, (2.3)] be employed as a lemma.
Lemma 2.4.
Let $\varphi:[0,\infty)\to[0,\infty)$ is of upper type $p_{1}\in(0,1]$ and lower type $p_{2}\in(0,1]$. Then, for ball $B(x_{0},r)$ and $t>0$,
$\|{\mathbf{1}}_{B(x_{0},\,tr)}\|_{L^{\varphi}({{{\mathbb{R}}}^{n}})}\leq\begin{cases}C^{+}t^{n/p_{1}}\|{\mathbf{1}}_{B(x_{0},\,r)}\|_{L^{\varphi}({{{\mathbb{R}}}^{n}})}&\text{as}\ t\in(0,1];\\ C^{-}t^{n/p_{2}}\|{\mathbf{1}}_{B(x_{0},\,r)}\|_{L^{\varphi}({{{\mathbb{R}}}^{n}})}&\text{as}\ t\in[1,\infty),\end{cases}$
where $C^{+}$ and $C^{-}$ are positive constants.
Lemma 2.5.
Let $0<\alpha<n$ and $\varphi$ be growth function with $n/(n+1)<i(\varphi)\leq 1$. Then $I_{\alpha}\in\mathcal{K_{\varphi,\,\alpha}}$.
Proof.
Accoording to[17, Theorem 3.1] and [9, Theorem 1.2.3], $I_{\alpha}$ has boundedness from $H^{1}(\mathbb{R}^{n})$ to $L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$, and $I_{\alpha}$ also bound from $L^{1}(\mathbb{R}^{n})$ to $L^{\frac{n}{n-\alpha},\,\infty}({{{\mathbb{R}}}^{n}})$. In order to prove this lemma, it suffices to show that $I_{\alpha}$ satisfies Definition 2.3(iii).
Now, we begin verify that $I_{\alpha}$ satisfies Definition 2.3(iii). From the condition $n/(n+1)<i(\varphi)$, we should choose $\sigma$ such that $\varphi$ is of lower type $\sigma$ with $n/(n+1)<\sigma<i(\varphi)$. So,
(2.7) $n\left(\frac{1}{\sigma}-1\right)<1.$
Suppose $b\in\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})$ and $a$ is $(\varphi,q,0)$ atom with $\operatorname{supp}a\subset B$. Denote
$p_{\alpha}=\frac{n}{n-\alpha}.$
Write
(2.8) $\|(b-b_{B})I_{\alpha}a\|_{L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})}$ $\leq\|(b-b_{B})\mathbf{1}_{2B}I_{\alpha}a\|_{L^{p_{\alpha}}({{{\mathbb{R}}}^{n}})}+\|(b-b_{B})\mathbf{1}_{(2B)^{c}}I_{\alpha}a\|_{L^{p_{\alpha}}({{{\mathbb{R}}}^{n}})}$
$=:{\rm I}+{\rm II}.$
To estimate $\rm I$}, if $q\in(1,n/\alpha)$ and $1/q_{0}=1/q-\alpha/n$, then we see that $I_{\alpha}$ is bounded from $L^{q}(\mathbb{R}^{n})$ to $L^{q_{0}}(\mathbb{R}^{n})$. Taking $r>1$ with $1/r+1/q_{0}=1/p_{\alpha}$, by Hölder inequality, Lemma 2.4 and [5, Lemma 3.4 and Lemma 5.5], it follows that
(2.9) ${\rm I}$ $\leq\left(\left(\int_{2B}\left|b(x)-b_{B}\right|^{p_{\alpha}\cdot\frac{r}{p_{\alpha}}}\,dx\right)^{\frac{p_{\alpha}}{r}}\cdot\left(\int_{2B}\left|I_{\alpha}a(x)\right|^{p_{\alpha}\cdot\frac{q_{0}}{p_{\alpha}}}\,dx\right)^{\frac{p_{\alpha}}{q_{0}}}\right)^{\frac{1}{p_{\alpha}}}$
$=\left(\int_{2B}|b(x)-b_{B}|^{r}\,dx\right)^{\frac{1}{r}}\left(\int_{2B}|I_{\alpha}a(x)|^{q_{0}}\,dx\right)^{\frac{1}{q_{0}}}$
$\lesssim\|a\|_{L^{q}(\mathbb{R}^{n})}\left(\left(\int_{2B}|b(x)-b_{2B}|^{r}\,dx\right)^{\frac{1}{r}}+\left(\int_{2B}|b_{2B}-b_{B}|^{r}\,dx\right)^{\frac{1}{r}}\right)$
$\lesssim\|a\|_{L^{q}(\mathbb{R}^{n})}\left(|2B|^{\frac{1}{r}-1}\|{\mathbf{1}}_{2B}\|_{L^{\varphi}(\mathbb{R}^{n})}\|b\|_{\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})}+|2B|^{\frac{1}{r}}2^{n/\sigma}|2B|^{-1}\|b\|_{\mathrm{BMO}_{\varphi}}\|{\mathbf{1}}_{B}\|_{L^{\varphi}(\mathbb{R}^{n})}\right)$
$\lesssim\|a\|_{L^{q}(\mathbb{R}^{n})}\ |B|^{\frac{1}{r}-1}\|{\mathbf{1}}_{B}\|_{L^{\varphi}(\mathbb{R}^{n})}\|b\|_{\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})}$
$=\|a\|_{L^{q}_{B}}|B|^{\frac{1}{q}+\frac{1}{r}-1}\|\mathbf{1}_{B}\|_{L^{\varphi}(\mathbb{R}^{n})}\|b\|_{\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})}$
$=\|a\|_{L^{q}_{B}}\|\mathbf{1}_{B}\|_{L^{\varphi}(\mathbb{R}^{n})}\|b\|_{\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})},$
where the penultimate step used $\|a\|_{L^{q}({{{\mathbb{R}}}^{n}})}=\|a\|_{L^{q}_{B}}|B|^{1/q}$ (see (2.6)) and the last step used
$\frac{1}{q}+\frac{1}{r}-1=\frac{1}{q_{0}}+\frac{\alpha}{n}+\frac{1}{p_{\alpha}}-\frac{1}{q_{0}}-1=\frac{\alpha}{n}+\frac{1}{p_{\alpha}}-1=0.$
Below, we consider $\rm II$}. Denote $B=B(x_{0},r)$. Due to the fact that $a$ is a $(\varphi,q,0)$-atom supported on a ball $B\subset{{{\mathbb{R}}}^{n}}$, we have $\operatorname{supp}a\subset B$ and $\int_{{{{\mathbb{R}}}^{n}}}a(x)\,dx=0$. For any $z\in B$ and $x\in(2^{j+1}B)\setminus(2^{j}B)$ with $j\in{\mathbb{N}}$, we find a point, which is labeled by $\xi_{z}$, lying on the straight segment joining $z$ to $x_{0}$, satisfying that
(2.10) $|I_{\alpha}a(x)|$ $=\left|\int_{B}a(z)\left(\frac{1}{|x-z|^{n-\alpha}}-\frac{1}{|x-x_{0}|^{n-\alpha}}\right)\,dz\right|$
$\leq\int_{B}\left|\frac{1}{|x-z|^{n-\alpha}}-\frac{1}{|x-x_{0}|^{n-\alpha}}\right||a(z)|\,dz$
$\leq\int_{B}|z-x_{0}|\left|(n-\alpha)|x-\xi_{z}|^{-(n-\alpha+1)}\right||a(z)|\,dz$
$\leq(n-\alpha)r\cdot\left(\frac{2^{j}r}{2}\right)^{-(n-\alpha+1)}\int_{B}|a(z)|\,dz,$
where the last step utilized
$|x-\xi_{z}|\geq|x-x_{0}|-|x_{0}-\xi_{z}|\geq|x-x_{0}|-\frac{|x-x_{0}|}{2}=\frac{|x-x_{0}|}{2}\geq\frac{2^{j}r}{2}.$
In addition, by $\|a\|_{L^{q}({{{\mathbb{R}}}^{n}})}=\|a\|_{L^{q}_{B}}|B|^{1-1/q^{\prime}}$, we get
$\int_{B}|a(z)|\,dz\leq|B|^{1/q^{\prime}}\|a\|_{L^{q}}=|B|\|a\|_{L^{q}_{B}},$
which and (2.10) imply
$|I_{\alpha}a(x)|\lesssim r\cdot r^{n}\cdot r^{-(n-\alpha+1)}\cdot 2^{-j(n-\alpha+1)}\|a\|_{L^{q}_{B}}=r^{\alpha}\cdot 2^{-j(n-\alpha+1)}\|a\|_{L^{q}_{B}}.$
This, along with lemma 2.4 and [5, Lemma 3.4 and Lemma 5.5], yields
${\rm II}$ $=\left(\int_{(2B)^{c}}\left|I_{\alpha}a(x)\left(b(x)-b_{B}\right)\right|^{p_{\alpha}}\,dx\right)^{\frac{1}{p_{\alpha}}}$
$\leq\sum_{j=1}^{\infty}\left(\int_{(2^{j+1}B)\setminus(2^{j}B)}\left|I_{\alpha}a(x)\left(b(x)-b_{B}\right)\right|^{p_{\alpha}}\,dx\right)^{\frac{1}{p_{\alpha}}}$
$\lesssim r^{\alpha}\|a\|_{L_{B}^{q}}\sum_{j=1}^{\infty}2^{-j(n-\alpha+1)}\Bigg[\left(\int_{2^{j+1}B}\left|b(x)-b_{2^{j+1}B}\right|^{p_{\alpha}}\,dx\right)^{\frac{1}{p_{\alpha}}}$
$\qquad+\left(\int_{2^{j+1}B}\left|b_{B}-b_{2^{j+1}B}\right|^{p_{\alpha}}\,dx\right)^{\frac{1}{p_{\alpha}}}\Bigg]$
$\lesssim r^{\alpha}\|a\|_{L_{B}^{q}}\sum_{j=1}^{\infty}2^{-j(n-\alpha+1)}\Big(|2^{j+1}B|^{\frac{1}{p_{\alpha}}-1}\|{\mathbf{1}}_{2^{j+1}B}\|_{L^{\varphi}({{{\mathbb{R}}}^{n}})}\|b\|_{\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})}$
$\qquad+|2^{j+1}B|^{\frac{1}{p_{\alpha}}}2^{\frac{(j+1)n}{\sigma}}|2^{j+1}B|^{-1}\|\mathbf{1}_{B}\|_{L^{\varphi}(\mathbb{R}^{n})}\|b\|_{\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})}\Big)$
$\lesssim r^{\alpha}\|a\|_{L^{q}_{B}}\sum_{j=1}^{\infty}2^{-j(n-\alpha+1)}|2^{j+1}B|^{\frac{1}{p_{\alpha}}-1}2^{\frac{(j+1)n}{\sigma}}\|\mathbf{1}_{B}\|_{L^{\varphi}({{{\mathbb{R}}}^{n}})}\|b\|_{\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})}$
$\approx r^{\alpha}r^{n\left(\frac{1}{p_{\alpha}}-1\right)}\|a\|_{L^{q}_{B}}\|\mathbf{1}_{B}\|_{L^{\varphi}({{{\mathbb{R}}}^{n}})}\|b\|_{\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})}\left(\sum_{j=1}^{\infty}2^{-j(1+n-\alpha)+n(j+1)(\frac{1}{p_{\alpha}}-1)+\frac{(j+1)n}{\sigma}}\right).$
Further, by $\alpha=-n\left(1/p_{\alpha}-1\right)$ and
$j(1+n-\alpha)+n(j+1)\left(\frac{1}{p_{\alpha}}-1\right)+\frac{(j+1)n}{\sigma}=-j\left(1-n\left(\frac{1}{\sigma}-1\right)\right)+\frac{n}{\sigma}-\alpha,$
we use (2.7) to obtain that
(2.11) $\mathrm{II}\lesssim\mathit{\|a\|_{L^{q}_{B}}\|\mathbf{1}_{B}\|_{L^{\varphi}({{{\mathbb{R}}}^{n}})}\|b\|_{\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})}}\sum_{j=1}^{\infty}2^{-j\left(1-n\left(\frac{1}{\sigma}-1\right)\right)}\lesssim\mathit{\|a\|_{L^{q}_{B}}\|\mathbf{1}_{B}\|_{L^{\varphi}({{{\mathbb{R}}}^{n}})}\|b\|_{\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})}}.$
Substituting (2.9) and (2.11) into (2.8) yields that $I_{\alpha}$ satisfies Definition 2.3(iii). ∎
3. Decomposition of commutator
This section is devoted to find a wavelet decomposition of commutator, where the wavelet basis is introduced in [18].
Denote $F=\{0,1\}^{n}\setminus\{(0,\dots,0)\}$, the function of all dyadic cubes in $\mathbb{R}^{n}$ is written as $\mathcal{D}$. Let $\{\psi_{I}^{\lambda}\}_{I\in\mathcal{D},\lambda\in F}$ be a normalized and orthogonal wavelet basis of $L^{2}(\mathbb{R}^{n})$ with compact support and sufficient smoothness (see [18]).
Any $g\in L^{2}(\mathbb{R}^{n})$ can be expanded as
$g=\sum_{I\in\mathcal{D}}\sum_{\lambda\in F}\langle g,\psi_{I}^{\lambda}\rangle\psi_{I}^{\lambda}\ \quad\text{for all}\ g\in L^{2}(\mathbb{R}^{n}).$
According to [1], we define the bilinear operators $\Pi_{i}\ (i=1,2,3,4)$ by
(3.12) $\Pi_{1}(g_{1},g_{2})=\sum_{\begin{subarray}{c}I,I^{\prime}\in\mathcal{D}\\ |I|=|I^{\prime}|\end{subarray}}\sum_{\lambda\in F}\langle g_{1},\varphi_{I}\rangle\langle g_{2},\psi_{I^{\prime}}^{\lambda}\rangle\varphi_{I}\psi_{I^{\prime}}^{\lambda},$
(3.13) $\Pi_{2}(g_{1},g_{2})=\sum_{\begin{subarray}{c}I,I^{\prime}\in\mathcal{D}\\ |I|=|I^{\prime}|\end{subarray}}\sum_{\lambda\in F}\langle g_{1},\psi_{I}^{\lambda}\rangle\langle g_{2},\varphi_{I^{\prime}}\rangle\psi_{I}^{\lambda}\varphi_{I^{\prime}},$
(3.14) $\Pi_{3}(g_{1},g_{2})=\sum_{\begin{subarray}{c}I,I^{\prime}\in\mathcal{D}\\ |I|=|I^{\prime}|\end{subarray}}\sum_{\begin{subarray}{c}\lambda,\lambda^{\prime}\in F\\ (I,\lambda)\neq(I^{\prime},\lambda^{\prime})\end{subarray}}\langle g_{1},\psi_{I}^{\lambda}\rangle\langle g_{2},\psi_{I^{\prime}}^{\lambda^{\prime}}\rangle\psi_{I}^{\lambda}\psi_{I^{\prime}}^{\lambda^{\prime}}$
and
(3.15) $\Pi_{4}(g_{1},g_{2})=\sum_{I\in\mathcal{D}}\sum_{\lambda\in F}\langle g_{1},\psi_{I}^{\lambda}\rangle\langle g_{2},\psi_{I}^{\lambda}\rangle(\psi_{I}^{\lambda})^{2}.$
Proposition 3.1.
Suppose that $\varphi$ is a growth function, its lower and upper critical indices satisfy (1.4). Respectively, each $\Pi_{i}$ ($i=1,2,3,4$) corresponds to the expression in (3.12), (3.13), (3.14) and (3.15). Consequently, the statements below are true.
  1. (i)
    Both $\Pi_{1}$ and $\Pi_{3}$ are bounded from $H^{\varphi}(\mathbb{R}^{n})\times\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})$ to $H^{1}(\mathbb{R}^{n})$;
  2. (ii)
    $\Pi_{4}$ is bounded from $H^{\varphi}(\mathbb{R}^{n})\times\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})$ to $L^{1}(\mathbb{R}^{n})$;
  3. (iii)
    Suppose that $I(\varphi)\in\left(n/(n+1),1\right)$ and (1.5) holds for $\varphi$, a sequence of dyadic cubes are denoted by $R_{j}$. Given $b\in\mathrm{BMO}_{\varphi}$ and a finite atomic decomposition $f=\sum_{j=1}^{M}a_{j}$ with $f\in H^{\varphi}({{{\mathbb{R}}}^{n}})$, where each $a_{j}$ can be written as a constant times a $(\varphi,\infty,s)$ atom, and $\operatorname{supp}a_{j}\subset R_{j}$. Then
    $\sum_{j=1}^{M}\big\|\Pi_{2}(a_{j},b-b_{B_{j}})\big\|_{H^{1}(\mathbb{R}^{n})}\;\lesssim\;\|f\|_{H^{\varphi}(\mathbb{R}^{n})}\,\|b\|_{\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})},$
    where $b_{B_{j}}$ is the average of $b$ over ball $B_{j}$ and $B_{j}\supseteq R_{j}$ is the smallest ball.
Proof.
Note that (i) and (ii) follow from [5, Propositions 3.14, 3.16 and 3.17]. By [5, (5.17)], we obtain that (iii) holds. ∎
Next we will establish a decomposition for the commutator $[b,T]$. The proof relies on a wavelet orthonormal basis $\{\psi_{l}^{\lambda}\}$ (see [18]) and the associated bilinear operators $\Pi_{i}$ introduced in [2].
Lemma 3.2.
Given a growth function $\varphi$ and its lower and upper critical indices satisfy (1.4) and (1.5). Let $\Pi_{4}$ be defined as (3.15) and $T\in K_{\varphi,\,\alpha}$ with $0<\alpha<n$. For any $b\in\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})$ and any $g\in H^{\varphi}(\mathbb{R}^{n})$, there exists an operator
$\mathfrak{T}:H^{\varphi}({{{\mathbb{R}}}^{n}})\times\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})\to L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$
such that
(3.16) $[b,T]g=\mathfrak{T}(g,\,b)-T(\Pi_{4}(g,\,b)).$
Proof.
Based on the argument of [16, Proposition 6.7], the following three steps will be used to prove the desired result.
Step 1: Verifying (3.16) under the case that $g$ has a finite linear combination of wavelets. We unitize [5, Lemma 3.3, (c) and (a)] to obtain that $g$ has finite atomic decomposition
$g=\sum^{M}_{k=1}a_{k}$
with
$a_{k}=\sum_{\begin{subarray}{c}I\in\mathcal{D}\\ I\subset R_{k}\end{subarray}}\sum_{\lambda\in F}\langle g,\psi^{\lambda}_{I}\rangle\psi^{\lambda}_{I}\qquad(k=1,\dots,M)$
and
$\Lambda_{\infty}\bigl(\{a_{k}\}_{k=1}^{M}\bigr)\lesssim\|g\|_{H^{\varphi}(\mathbb{R}^{n})},$
where $\operatorname{supp}a_{k}\subseteq R_{k}$ and $\langle g,\psi^{\lambda}_{I}\rangle$ is the coefficients of $a_{k}$. Further, let $B_{k}$ be the smallest ball containing dyadic cube $R_{k}$. According to [5, Proposition 2.5], we have
(3.17) $\sum_{k=1}^{M}\|a_{k}\|_{L^{\infty}_{B_{k}}}\|\mathbf{1}_{B_{k}}\|_{L^{\varphi}(\mathbb{R}^{n})}\lesssim\Lambda_{\infty}\bigl(\{a_{k}\}_{k=1}^{M}\bigr)\lesssim\|g\|_{H^{\varphi}(\mathbb{R}^{n})}.$
Denote
$\mathcal{W}(g,\,b):=T\left(\sum_{k=1}^{M}\left[(b-b_{B_{k}})a_{k}-\Pi_{2}\left(a_{k},b-b_{B_{k}}\right)\right]\right).$
For almost every $x\in\mathbb{R}^{n}$, using a standard algebraic identity (see [16, (6.9)]), it holds
$b(x)g(\cdot)-b(\cdot)g(\cdot)=$ $\sum_{k=1}^{M}\Bigl[\bigl(b(x)-b_{B_{k}}\bigr)a_{k}(\cdot)-\Pi_{2}\bigl(a_{k},b-b_{B_{k}}\bigr)(\cdot)\Bigr]$
$\quad-\Pi_{1}(g,b)(\cdot)-\Pi_{3}(g,b)(\cdot)-\Pi_{4}(g,b)(\cdot).$
This implies
(3.18) $[b,T]g=\mathcal{W}(g,\,b)-T(\Pi_{1}(g,\,b))-T(\Pi_{3}(g,\,b))-T(\Pi_{4}(g,\,b)).$
Now, if we define
$\mathfrak{T}(g,\,b)=\mathcal{W}(g,\,b)-T(\Pi_{1}(g,\,b))-T(\Pi_{3}(g,\,b)),$
then immediately obtain
$b,T]g=\mathfrak{T}(g,\,b)-T(\Pi_{4}(g,\,b)),$
which is the desired equality.
Step 2: For $b\in\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})$ and $g\in H^{\varphi}({{{\mathbb{R}}}^{n}})$ with finite linear combination of wavelets, it follows that
$\|\mathfrak{T}(g,\,b)\|_{L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})}\lesssim\|g\|_{H^{\varphi}({{{\mathbb{R}}}^{n}})}\|b\|_{\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})}.$
Using Proposition 3.1, for arbitrary $g\in H^{\varphi}({{{\mathbb{R}}}^{n}})$ and $b\in\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})$,
$\|\Pi_{i}(g,\,b)\|_{H^{1}({{{\mathbb{R}}}^{n}})}\lesssim\|g\|_{H^{\varphi}({{{\mathbb{R}}}^{n}})}\|b\|_{\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})},\quad i=1,3.$
By this, $T\in K_{\varphi,\,\alpha}$, Proposition 3.1(iii) , (3.17) and $\|a_{k}\|_{L^{q}_{B_{k}}}\leq\|a_{k}\|_{L^{\infty}_{B_{k}}}$ with $q\in(1,n/\alpha)$, we see that
$\|\mathfrak{T}(g,b)\|_{L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})}$ $\leq\|\mathcal{U}(g,b)\|_{L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})}+\|T(\Pi_{1}(g,b))\|_{L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})}+\|T(\Pi_{3}(g,b))\|_{L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})}$
$\lesssim\sum_{k=1}^{M}\left\|(b(\cdot)-b_{B_{k}})Ta_{k}\right\|_{L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})}+\left\|T\left(\sum_{k=1}^{M}\Pi_{2}(a_{k},b-b_{B_{k}})\right)\right\|_{L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})}$
$\qquad+\|\Pi_{1}(g,b)\|_{H^{1}(\mathbb{R}^{n})}+\|\Pi_{3}(g,b)\|_{H^{1}(\mathbb{R}^{n})}$
$\lesssim\|g\|_{H^{\varphi}(\mathbb{R}^{n})}\|b\|_{\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})}+\sum_{k=1}^{M}\|a_{k}\|_{L^{q}_{B_{k}}}\|\mathbf{1}_{B_{k}}\|_{L^{\varphi}(\mathbb{R}^{n})}\|b\|_{\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})}$
$\qquad+\sum_{k=1}^{M}\left\|\Pi_{2}(a_{k},b-b_{B_{k}})\right\|_{H^{1}(\mathbb{R}^{n})}$
$\lesssim\|g\|_{H^{\varphi}(\mathbb{R}^{n})}\|b\|_{\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})},$
as desired.
Step 3: The proof of (3.16). Based on the facts obtained in Steps 1 and 2, we repeat the argument of [5, (5.18)] to obtain (3.16). This implies Lemma 3.2. ∎
4. Proofs of Theorems 1.5 and 1.6
The primary objective of this section is to complete the proofs of Theorems 1.5 and 1.6 by using Lemma 4.1. We first recall the following lemma from [5, Theorem 5.10].
Lemma 4.1.
Suppose that $\varphi$ is growth function and it satisfies (1.4) and (1.5). If $b\in\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})$, then $g\in H^{\varphi}_{b}(\mathbb{R}^{n})$ is equivalent to $\Pi_{4}(g,\,b)\in H^{1}(\mathbb{R}^{n})$ with
$\|g\|_{H^{\varphi}_{b}(\mathbb{R}^{n})}\approx\|g\|_{H^{\varphi}(\mathbb{R}^{n})}\|b\|_{\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})}+\left\|\Pi_{4}(g,\,b)\right\|_{H^{1}(\mathbb{R}^{n})}.$
Proof of Theorem 1.5.
Under the assumptions that $b\in\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})$ and $g\in H^{\varphi}_{b}(\mathbb{R}^{n})$, by Lemma 4.1, we have $\Pi_{4}(g,\,b)\in H^{1}(\mathbb{R}^{n})$ and
(4.19) $\|g\|_{H^{\varphi}_{b}(\mathbb{R}^{n})}\approx\|g\|_{H^{\varphi}(\mathbb{R}^{n})}\|b\|_{\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})}+\|\Pi_{4}(g,\,b)\|_{H^{1}(\mathbb{R}^{n})}.$
Based on Lemma 2.5, we use Lemma 3.2 to obtain that there exists operator
$\mathfrak{T}:H^{\varphi}(\mathbb{R}^{n})\times\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})\to L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$
such that
$b,I_{\alpha}]g=\mathfrak{T}(g,\,b)-{I_{\alpha}}(\Pi_{4}(g,\,b)).$
From this and (1.2), it follows that
$\|[b,I_{\alpha}]g\|_{L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})}$ $=\|\mathfrak{T}(g,\,b)-I_{\alpha}(\Pi_{4}(g,\,b))\|_{L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})}$
$\leq\|\mathfrak{T}(g,\,b)\|_{L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})}+\|I_{\alpha}(\Pi_{4}(g,\,b))\|_{L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})}$
$\lesssim\|g\|_{H^{\varphi}(\mathbb{R}^{n})}\|b\|_{\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})}+\|\Pi_{4}(g,\,b)\|_{H^{1}(\mathbb{R}^{n})}$
$\approx\|g\|_{H^{\varphi}_{b}(\mathbb{R}^{n})},$
which proves Theorem1.5. ∎
Proof of Theorem 1.6.
By Lemma 3.2, we find that there exists operator
$\mathfrak{T}:H^{\varphi}(\mathbb{R}^{n})\times\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})\to L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$
such that for any $b\in\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})$ and $g\in H^{\varphi}(\mathbb{R}^{n})$,
$b,I_{\alpha}]g(y)=\mathfrak{T}(g,\,b)(y)-{I_{\alpha}}(\Pi_{4}(g,\,b))(y)\ \quad\text{for a. e.}\,y\in{{{\mathbb{R}}}^{n}}.$
So, using (1.1) and Proposition 3.1, we have
$\|[b,I_{\alpha}]g\|_{L^{\frac{n}{n-\alpha},\,{\infty}}(\mathbb{R}^{n})}$ $\lesssim\|\mathfrak{T}(g,\,b)\|_{L^{\frac{n}{n-\alpha},\,{\infty}}(\mathbb{R}^{n})}+\|I_{\alpha}(\Pi_{4}(g,\,b))\|_{L^{\frac{n}{n-\alpha},\,{\infty}}(\mathbb{R}^{n})}$
$\lesssim\|\mathfrak{T}(g,\,b)\|_{L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})}+\|\Pi_{4}(g,\,b)\|_{L^{1}({{{\mathbb{R}}}^{n}})}$
$\lesssim\|g\|_{H^{\varphi}(\mathbb{R}^{n})}\|b\|_{\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})}.$
This implies Theorem1.5. ∎

### 28 Apr 2026 v2

@@
 $I_{\alpha}(g)(x):=\int_{\mathbb{R}^{n}}\frac{g(z)}{\left|x-z\right|^{n-\alpha}}\,dz\quad\text{for all}\,x\in{{{\mathbb{R}}}^{n}}.$
-Two classical results (see [9, Theorem 1.2.3] and [11, Theorem 4]) asserted that there exists constants $C$ such that
+Two classical results (see [10, Theorem 1.2.3] and [12, Theorem 4]) asserted that there exists constants $C$ such that
 (1.1) $\|I_{\alpha}g\|_{L^{\frac{n}{n-\alpha},\,{\infty}}(\mathbb{R}^{n})}\leq C\|g\|_{L^{1}({{{\mathbb{R}}}^{n}})}$
@@
 with $1<p<q<\infty$ and $1/q-1/p=\alpha/n$.
-In the framework of Hardy spaces, the boundedness of $I_{\alpha}$ is considered by Stein and Weiss [21, Theorem G] in their pioneering work on multi-variable harmonic functions, they established that
+In the framework of Hardy spaces, the boundedness of $I_{\alpha}$ is considered by Stein and Weiss [26, Theorem G] in their pioneering work on multi-variable harmonic functions, they established that
 (1.2) $\|I_{\alpha}g\|_{L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})}\leq C\|g\|_{H^{1}({{{\mathbb{R}}}^{n}})},$
@@
 A famous theorem of Chanillo [4] characterizes the $\mathrm{BMO}({{{\mathbb{R}}}^{n}})$ space via the boundedness of $b,I_{\alpha}]$, that is, $b\in\mathrm{BMO}({{{\mathbb{R}}}^{n}})$ if and only if the commutator $b,I_{\alpha}]$ is bounded from $L^{p}({{{\mathbb{R}}}^{n}})$ to $L^{q}({{{\mathbb{R}}}^{n}})$, where $0<\alpha<n$ and $1<p<\frac{n}{\alpha}$ with $1/q=1/p-\alpha/n$.
-Paluszyński [19, Theorem 1.1] proved that $b,I_{\alpha}]$ does not map from $H^{1}({{{\mathbb{R}}}^{n}})$ into $L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$ continuously.
+Paluszyński [22, Theorem 1.1] proved that $b,I_{\alpha}]$ does not map from $H^{1}({{{\mathbb{R}}}^{n}})$ into $L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$ continuously.
 Later, Pérez [23] introduced a subspace $\mathcal{H}^{1}_{b}({{{\mathbb{R}}}^{n}})\subset H^{1}({{{\mathbb{R}}}^{n}})$ in the study of commutator of Calderón-Zygmund operator.
@@
 Moreover, Ky [14] also proved that the $b,I_{\alpha}]$ maps $H^{1}({{{\mathbb{R}}}^{n}})$ continuously into the weak Lebesgue space $L^{\frac{n}{n-\alpha},\,\infty}({{{\mathbb{R}}}^{n}})$.
-For more related results, we refer to the interested reader to [3, 4, 10, 15, 20].
+For more related results, we refer to the interested reader to [1, 5, 11, 13, 16, 21, 24, 25].
 Motivated by the above works, the present paper studies the boundedness of commutator generated by Orlicz-$\mathrm{BMO}$ function and fractional integral operator within the framework of Orlicz-type spaces.
@@
 $g_{B}:=\frac{1}{|B|}\int_{B}g(x)\,dx.$
-If $\varphi(t)=t$, $\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})=\mathrm{BMO}(\mathbb{R}^{n})$. For more research on $\mathrm{BMO}$ type spaces, see [6, 7, 8, 22].
+If $\varphi(t)=t$, $\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})=\mathrm{BMO}(\mathbb{R}^{n})$. For more research on $\mathrm{BMO}$ type spaces, see [1, 7, 8, 9, 27].
 The Orlicz-Lebesgue space $L^{\varphi}({{{\mathbb{R}}}^{n}})$ contains all locally integrable functions $g$ for which
@@
 Remark 1.7.
-Theorems 1.5 and 1.6 are a further extension for the results of Ky [13, Corollary 8.1 and Theorem 8.2] in the Orlicz-type spaces framework. A study related to the Calderón-Zygmund operator, see [5].
+Theorems 1.5 and 1.6 are a further extension for the results of Ky [14, Corollary 8.1 and Theorem 8.2] in the Orlicz-type spaces framework. A study related to the Calderón-Zygmund operator, see [6].
 In Section 2, we introduce the operator class $\mathcal{K}_{\varphi,\,\alpha}$ and prove $I_{\alpha}\in\mathcal{K}_{\varphi,\,\alpha}$.
@@
     $\|(b-b_{B})Ra\|_{L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})}\leq C\|a\|_{L^{q}_{B}}\|{\mathbf{1}_{B}\|_{L^{\varphi}({{{\mathbb{R}}}^{n}})}\|b\|_{\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})}}.$
-To facilitate subsequent proof, it is proposed that the result in [5, (2.3)] be employed as a lemma.
+To facilitate subsequent proof, it is proposed that the result in [6, (2.3)] be employed as a lemma.
 Lemma 2.4.
@@
 Proof.
-Accoording to[17, Theorem 3.1] and [9, Theorem 1.2.3], $I_{\alpha}$ has boundedness from $H^{1}(\mathbb{R}^{n})$ to $L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$, and $I_{\alpha}$ also bound from $L^{1}(\mathbb{R}^{n})$ to $L^{\frac{n}{n-\alpha},\,\infty}({{{\mathbb{R}}}^{n}})$.
+Accoording to[19, Theorem 3.1] and [10, Theorem 1.2.3], $I_{\alpha}$ has boundedness from $H^{1}(\mathbb{R}^{n})$ to $L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$, and $I_{\alpha}$ also bounded from $L^{1}(\mathbb{R}^{n})$ to $L^{\frac{n}{n-\alpha},\,\infty}({{{\mathbb{R}}}^{n}})$.
 In order to prove this lemma, it suffices to show that $I_{\alpha}$ satisfies Definition 2.3(iii).
@@
 To estimate $\rm I$}, if $q\in(1,n/\alpha)$ and $1/q_{0}=1/q-\alpha/n$, then we see that $I_{\alpha}$ is bounded from $L^{q}(\mathbb{R}^{n})$ to $L^{q_{0}}(\mathbb{R}^{n})$.
-Taking $r>1$ with $1/r+1/q_{0}=1/p_{\alpha}$, by Hölder inequality, Lemma 2.4 and [5, Lemma 3.4 and Lemma 5.5], it follows that
+Taking $r>1$ with $1/r+1/q_{0}=1/p_{\alpha}$, by Hölder inequality, Lemma 2.4 and [6, Lemma 3.4 and Lemma 5.5], it follows that
 (2.9) ${\rm I}$ $\leq\left(\left(\int_{2B}\left|b(x)-b_{B}\right|^{p_{\alpha}\cdot\frac{r}{p_{\alpha}}}\,dx\right)^{\frac{p_{\alpha}}{r}}\cdot\left(\int_{2B}\left|I_{\alpha}a(x)\right|^{p_{\alpha}\cdot\frac{q_{0}}{p_{\alpha}}}\,dx\right)^{\frac{p_{\alpha}}{q_{0}}}\right)^{\frac{1}{p_{\alpha}}}$
@@
 $|I_{\alpha}a(x)|\lesssim r\cdot r^{n}\cdot r^{-(n-\alpha+1)}\cdot 2^{-j(n-\alpha+1)}\|a\|_{L^{q}_{B}}=r^{\alpha}\cdot 2^{-j(n-\alpha+1)}\|a\|_{L^{q}_{B}}.$
-This, along with lemma 2.4 and [5, Lemma 3.4 and Lemma 5.5], yields
+This, along with lemma 2.4 and [6, Lemma 3.4 and Lemma 5.5], yields
 ${\rm II}$ $=\left(\int_{(2B)^{c}}\left|I_{\alpha}a(x)\left(b(x)-b_{B}\right)\right|^{p_{\alpha}}\,dx\right)^{\frac{1}{p_{\alpha}}}$
@@
 Proof.
-Note that (i) and (ii) follow from [5, Propositions 3.14, 3.16 and 3.17]. By [5, (5.17)], we obtain that (iii) holds. ∎
+Note that (i) and (ii) follow from [6, Propositions 3.14, 3.16 and 3.17]. By [6, (5.17)], we obtain that (iii) holds. ∎
 Next we will establish a decomposition for the commutator $[b,T]$.
@@
 Given a growth function $\varphi$ and its lower and upper critical indices satisfy (1.4) and (1.5).
-Let $\Pi_{4}$ be defined as (3.15) and $T\in K_{\varphi,\,\alpha}$ with $0<\alpha<n$.
+Let $\Pi_{4}$ be defined as (3.15) and $T\in\mathcal{K}_{\varphi,\,\alpha}$ with $0<\alpha<n$.
 For any $b\in\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})$ and any $g\in H^{\varphi}(\mathbb{R}^{n})$, there exists an operator
@@
 Proof.
-Based on the argument of [16, Proposition 6.7], the following three steps will be used to prove the desired result.
-Step 1: Verifying (3.16) under the case that $g$ has a finite linear combination of wavelets. We unitize [5, Lemma 3.3, (c) and (a)] to obtain that $g$ has finite atomic decomposition
+Based on the argument of [18, Proposition 6.7], the following three steps will be used to prove the desired result.
+Step 1: Verifying (3.16) under the case that $g$ has a finite linear combination of wavelets. We unitize [6, Lemma 3.3, (c) and (a)] to obtain that $g$ has finite atomic decomposition
 $g=\sum^{M}_{k=1}a_{k}$
@@
 Further, let $B_{k}$ be the smallest ball containing dyadic cube $R_{k}$.
-According to [5, Proposition 2.5], we have
+According to [6, Proposition 2.5], we have
 (3.17) $\sum_{k=1}^{M}\|a_{k}\|_{L^{\infty}_{B_{k}}}\|\mathbf{1}_{B_{k}}\|_{L^{\varphi}(\mathbb{R}^{n})}\lesssim\Lambda_{\infty}\bigl(\{a_{k}\}_{k=1}^{M}\bigr)\lesssim\|g\|_{H^{\varphi}(\mathbb{R}^{n})}.$
@@
 $\mathcal{W}(g,\,b):=T\left(\sum_{k=1}^{M}\left[(b-b_{B_{k}})a_{k}-\Pi_{2}\left(a_{k},b-b_{B_{k}}\right)\right]\right).$
-For almost every $x\in\mathbb{R}^{n}$, using a standard algebraic identity (see [16, (6.9)]), it holds
+For almost every $x\in\mathbb{R}^{n}$, using a standard algebraic identity (see [18, (6.9)]), it holds
 $b(x)g(\cdot)-b(\cdot)g(\cdot)=$ $\sum_{k=1}^{M}\Bigl[\bigl(b(x)-b_{B_{k}}\bigr)a_{k}(\cdot)-\Pi_{2}\bigl(a_{k},b-b_{B_{k}}\bigr)(\cdot)\Bigr]$
@@
 $\|\Pi_{i}(g,\,b)\|_{H^{1}({{{\mathbb{R}}}^{n}})}\lesssim\|g\|_{H^{\varphi}({{{\mathbb{R}}}^{n}})}\|b\|_{\mathrm{BMO}_{\varphi}({{{\mathbb{R}}}^{n}})},\quad i=1,3.$
-By this, $T\in K_{\varphi,\,\alpha}$, Proposition 3.1(iii) , (3.17) and $\|a_{k}\|_{L^{q}_{B_{k}}}\leq\|a_{k}\|_{L^{\infty}_{B_{k}}}$ with $q\in(1,n/\alpha)$, we see that
+By this, $T\in\mathcal{K}_{\varphi,\,\alpha}$, Proposition 3.1(iii) , (3.17) and $\|a_{k}\|_{L^{q}_{B_{k}}}\leq\|a_{k}\|_{L^{\infty}_{B_{k}}}$ with $q\in(1,n/\alpha)$, we see that
 $\|\mathfrak{T}(g,b)\|_{L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})}$ $\leq\|\mathcal{U}(g,b)\|_{L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})}+\|T(\Pi_{1}(g,b))\|_{L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})}+\|T(\Pi_{3}(g,b))\|_{L^{\frac{n}{n-\alpha}}(\mathbb{R}^{n})}$
@@
 as desired.
-Step 3: The proof of (3.16). Based on the facts obtained in Steps 1 and 2, we repeat the argument of [5, (5.18)] to obtain (3.16). This implies Lemma 3.2. ∎
+Step 3: The proof of (3.16). Based on the facts obtained in Steps 1 and 2, we repeat the argument of [6, (5.18)] to obtain (3.16). This implies Lemma 3.2. ∎
 4. Proofs of Theorems 1.5 and 1.6
-The primary objective of this section is to complete the proofs of Theorems 1.5 and 1.6 by using Lemma 4.1. We first recall the following lemma from [5, Theorem 5.10].
+The primary objective of this section is to complete the proofs of Theorems 1.5 and 1.6 by using Lemma 4.1. We first recall the following lemma from [6, Theorem 5.10].
 Lemma 4.1.